Hi
Mark went to buy an equal number of apples and oranges. The apples were 5 for $3 and the cost of 2 oranges was $3,30. He paid $31.50 more for the oranges than for the apples. How many oranges and apples did he buy altogether.
Thanks

Answer by ikleyn(52878)

.
Mark went to buy an equal number of apples and oranges.
The apples were 5 for $3 and the cost of 2 oranges was $3.30.
He paid $31.50 more for the oranges than for the apples.
How many oranges and apples did he buy altogether.
~~~~~~~~~~~~~~~~

The apples were 5 for $3.00,  or $0.60 each.

The oranges were 2 for $3.30, or $1.65 each.


x oranges and x apples (an equal number).


0.60x dollars for x apples and 1.65x dollars for x oranges.


The difference in cost was 31.50 dollars

    1.65x - 0.60x = 31.50 dollars.


It is your basic equation.
To solve it, simplify

      1.05x = 31.50

          x = 31.50/1.05 = 30.


ANSWER.  30 apples and 30 oranges, or 30+30 = 60 fruits altogether.

Solved.